# The Degree of (Extended) Justified Representation and Its Optimization

### v1

@@ Abstract
Justified Representation (JR)/Extended Justified Representation (EJR) is a desirable axiom in multiwinner approval voting. In contrast to (E)JR only requires at least one voter to be represented in every cohesive group, we study its optimization version that maximizes the number of represented voters in each group. Given an instance, we say a winning committee provides an (E)JR degree of $c$ if at least $c$ voters in each $\ell$-cohesive group have approved $\ell$ winning candidates. Hence, every (E)JR committee provides the (E)JR degree of at least $1$. Besides proposing this new property, we propose the optimization problem of finding a winning committee that achieves the maximum possible (E)JR degree, called MDJR and MDEJR, corresponding to JR and EJR respectively.
We study the computational complexity and approximability of MDJR of MDEJR. An (E)JR committee, which can be found in polynomial time, straightforwardly gives a $(k/n)$-approximation. On the other hand, we show that it is NP-hard to approximate MDJR and MDEJR to within a factor of $\left(k/n\right)^{1-\epsilon}$, for any $\epsilon>0$, which complements the approximation. Next, we study the fixed-parameter-tractability of this problem. We show that both problems are W[2]-hard if $k$, the size of the winning committee, is specified as the parameter. However, when $c_{\text{max}}$, the maximum value of $c$ such that a committee that provides an (E)JR degree of $c$ exists, is additionally given as a parameter, we show that both MDJR and MDEJR are fixed-parameter-tractable.
@@ 1 Introduction
[2] Given a ballot profile $\mathbf{A}=(A_{1},\dots,A_{n})$ over a candidate set $C$ and a committee size $k$, we say that a set of candidates $W$ of size $|W|=k$ satisfies EJR for $(\mathbf{A},k)$ if, for every $\ell$-cohesive group with every $\ell\in[k]$, there are at least one voter which approves at least $\ell$ candidates in $W$.
@@
While there is a wealth of concepts associated with JR, there are instances where comparing the performance of different winning committees becomes impossible. To illustrate this situation, consider a brief example.
@@
  • each of voters $c_{4},c_{5},c_{6}$ is approved by voters $1,2,4,5,7,8$.
@@ 1.2 Our Technical Contributions
  • We show that it is NP-hard to approximate MDJR (MDEJR) within a factor of $(k/n)^{1-\epsilon}$ for any $\epsilon>0$. This complements the result that an EJR committee (i.e., winner committee with both JR degree and EJR degree at least $1$) always exists and can be computed in polynomial time, which implies a $(k/n)$-approximation for both MDJR and MDEJR.
@@ 2 Inapproximability of MDJR and MDEJR
We have seen in the previous section that an EJR winner committee always exists and can be computed in polynomial time. Since an EJR winner committee achieves a degree of at least $1$ for both JR and EJR and the maximum possible degree is $n/k$, this implies a $(k/n)$-approximation algorithm for both MDJR and MDEJR problems. To complement this positive result, we will show that both MDJR and MDEJR are NP-hard to approximate to within a factor of $\left(\frac{k}{n}\right)^{1-\epsilon}$ for any $\epsilon>0$.
@@
We will simultaneously prove the two theorems by constructing a hard ballot instance that is used for proving both theorems. We will make sure the instance we constructed has no $\ell$-cohesive group with $\ell>1$. Notice that the JR degree and the EJR degree for any committee are always the same for instances with only $1$-cohesive groups.
2.1 Proof of Theorem 1 and 2
Before we prove the theorem, we first introduce a NP-hard problem: sparse-SAT problem. One can find variations of SAT problem and a similar argument in [36].
@@ 2.1 Proof of Theorem 1 and 2
Given a CNF formula $\phi$ that, for any variables $x$ and $y$, at most one clause contains both $x$ (or $\neg x$) and $y$ (or $\neg y$), decide if there is a value assignment to the variables to make $\phi$ true.
To see the NP-hardness, it can be reduced from the SAT problem. Start with any SAT instance. Without loss of generality, suppose each variable, $x$ or $\neg x$, appears in each clause at most once. For each variable $x$ that there are more than two clauses containing $x$ or $\neg x$ perform the following procedure: Suppose $x$ appears in $k$ clauses. Create $k$ new variables $x_{1},\dots,x_{k}$ and replace the $i$th occurrence of $x$ with $x_{i}$ (and $\neg x$ is replaced by $\neg x_{i}$, respectively), $i=1,\dots,k$. Append the clause $(x_{i}\vee\neg x_{i+1})$ for $i=1,\dots,k-1$ and the clause $(x_{k}\vee\neg x_{1})$. Note that, in the new instance, variable $x_{i}$ and $y_{j}$ appear in a clause only when the $i$th occurrence of $x$ and the $j$th occurrence of $y$ in $\phi$ are in the same clause, so the new instance satisfies the limitation of sparse-SAT problem.
@@
Proof of Theorem 1 and 2.
@@
We reduce from sparse-SAT problem. Given any sparse-SAT instance $\phi$, suppose there are $\bar{n}$ clauses and $\bar{m}$ different variables (say $x_{1},\dots,x_{\bar{m}}$). We consider a ABC voting instance with $13\bar{m}+\bar{n}+$ 1 candidates, $c_{1},\dots,c_{2\bar{m}}$, $s_{1},\dots,s_{\bar{n}}$, $t_{1},\dots,t_{\bar{m}}$, $d$ and $\bar{n}\bar{m}+\bar{m}^{2}+\bar{n}+\bar{m}$ voters. We want to select a committee of size $\bar{m}+1$. Hence, we care about the cohesive group of size $\frac{\bar{n}\bar{m}+\bar{m}^{2}+\bar{n}+\bar{m}}{\bar{m}+1}=\bar{n}+\bar{m}$. First, for each variable $x_{j}$ and its complement $\neg x_{j}$, we create two corresponding candidates $c_{2j-1}$ and $c_{2j}$ and a group $T_{j}$ of $\bar{m}$ voters who approve $c_{2j-1}$ and $c_{2j}$. For the $i$th clause, we create a group $S_{i}$ of $\bar{m}$ voters. All voters in group $S_{i}$ approve $c_{2j-1}$ if $x_{j}$ occurs in the $i$th clause and $c_{2j}$ if $\neg x_{j}$ occurs in the $i$th clause. All voters in group $S_{i}$ approve $s_{i}$ and voters in group $T_{i}$ approve $t_{i}$ additionally. We create a set $D$ of $\bar{n}$ voters who approve $s_{1},\dots,s_{\bar{n}},t_{1},\dots,t_{\bar{m}},d$. Hence, for each $i\in[\bar{n}]$, $S_{i}\cup D$ forms a $1$-cohesive group and for each $j\in[\bar{m}]$, $T_{j}\cup D$ forms a $1$-cohesive group. Moreover, we create a set $D^{+}$ of $\bar{m}$ voters who approve $d$. Hence, $D\cup D^{+}$ forms a $1$-cohesive group.
@@
Notably, there is no $2$-cohesive group. First, candidate $s_{i}$ for $i\in[\bar{n}]$ or $t_{j}$ for $j\in[\bar{m}]$ or $d$ has only $\bar{n}+\bar{m}$ voters approving them as constructed above. For candidates $c_{1},\dots,c_{2\bar{m}}$, we will show that no $2(\bar{n}+\bar{m})$ voters have two common approved candidates. Since in a sparse-SAT instance, for any two variables $x_{i}$ and $x_{j}$, at most one clause contains them or their complement, for any two candidates $c_{2i}$ or $c_{2i-1}$ and $c_{2j}$ or $c_{2j-1}$, at most one group $S_{k}$ of $\bar{m}$ voters approves them simultaneously. Moreover, only $\bar{m}$ voters in group $T_{i}$ approve both $c_{2i-1}$ and $c_{2i}$. So there is no $2$-cohesive group or $\ell$-cohesive group for $l>2$. In this case, the EJR degree equals to JR degree.
@@
If there is a value assignment to the variables to make $\phi$ true, then MDEJR will achieve the EJR degree of $\bar{n}+\bar{m}$. To see this, we elect $c_{2j-1}$ as a winner if $x_{j}$ assigned true or we elect $c_{2j}$ for all $j\in[\bar{m}]$. Then we elect $d$ as the winner. We can verify that all voters in $S_{i}$ and $T_{j}$ are covered for every $i\in[\bar{n}],j\in[\bar{m}]$. In addition, $d$ covers all voters in $D$ and $D^{+}$. Hence, all voters are covered and the JR degree equals the size of $1$-cohesive group, $\bar{n}+\bar{m}$, and so does the EJR degree.
@@
If there does not exist a value assignment to the variables to make $\phi$ true, then MDEJR will achieve the EJR degree of at most $\bar{n}$. To see this, we prove it by contradiction. Assume that MDEJR can achieve the EJR degree larger than $\bar{n}$. If $d$ is not elected, no voter in $D^{+}$ can be covered since they only approve $d$. Hence, for cohesive group $D\cup D^{+}$, there are at most $\bar{n}$ voters can be covered, which makes a contradiction. If $d$ is elected, there will be $\bar{m}$ vacant seats. If voters in $T_{j}$ are not covered for $j\in[\bar{m}]$, there are at most $\bar{n}$ voters covered in cohesive group $D\cup T_{j}$, a contradiction, so all voters in group $T_{j}$ must be covered for all $j\in[\bar{m}]$, indicating that exact one of $3$ candidates $\{c_{2j-1},c_{2j},t_{j}\}$ is elected. Next, we will show that there is some group among $S_{1},\dots,S_{\bar{n}}$ that cannot be covered. To see this, suppose there is a committee $W^{\prime}$ of size $\bar{m}$ which covers all groups $S_{1},\dots,S_{\bar{n}}$. For every $t_{j}\in W^{\prime}$, we can find $c_{2j-1}$ or $c_{2j}$ to replace $t_{j}$ since $t_{j}$ only covers group $T_{j}$. Hence, we can find one candidate in $\{c_{2i-1},c_{2i}\}$ for all $i\in[\bar{m}]$ to cover all groups, implying that there exists a value assignment to the variables ($x_{i}$ is assigned true if $c_{2i-1}$ is elected or false otherwise) to make $\phi$ true, leading to a contradiction. Therefore, there is some group among $S_{1},\dots,S_{\bar{n}}$ that cannot be covered. Without loss of generality, we assume that it is $S_{i}$. Then, for cohesive group $S_{i}\cup D$, there are at most $\bar{n}$ voters that can be covered, which makes a contradiction.

### v2

@@ Abstract
Justified Representation (JR)/Extended Justified Representation (EJR) is a desirable axiom in multiwinner approval voting. In contrast to that (E)JR only requires at least one voter to be represented in every cohesive group, we study its optimization version that maximizes the number of represented voters in each group. Given an instance, we say a winning committee provides a JR degree (EJR degree, resp.) of $c$ if at least $c$ voters in each $\ell$-cohesive group ($1$-cohesive group, resp.) have approved $\ell$ ($1$, resp.) winning candidates. Hence, every (E)JR committee provides the (E)JR degree of at least $1$. Besides proposing this new property, we propose the optimization problem of finding a winning committee that achieves the maximum possible (E)JR degree, called MDJR and MDEJR, corresponding to JR and EJR respectively.
We study the computational complexity and approximability of MDJR of MDEJR. An (E)JR committee, which can be found in polynomial time, straightforwardly gives a $(k/n)$-approximation. We also show that the original algorithms for finding a JR and an EJR winner committee are also $1/k$ and $1/(k+1)$ approximation algorithms for MDJR and MDEJR respectively. On the other hand, we show that it is NP-hard to approximate MDJR and MDEJR to within a factor of $\left(k/n\right)^{1-\epsilon}$ and to within a factor of $(1/k)^{1-\varepsilon}$, for any $\epsilon>0$, which complements the positive results. Next, we study the fixed-parameter-tractability of this problem. We show that both problems are W[2]-hard if $k$, the size of the winning committee, is specified as the parameter. However, when $c_{\text{max}}$, the maximum value of $c$ such that a committee that provides an (E)JR degree of $c$ exists, is additionally given as a parameter, we show that both MDJR and MDEJR are fixed-parameter-tractable.
@@ 1 Introduction
[2] Given a ballot profile $\mathbf{A}=(A_{1},\dots,A_{n})$ over a candidate set $C$ and a committee size $k$, we say that a set of candidates $W$ of size $|W|=k$ satisfies EJR for $(\mathbf{A},k)$ if, for every $\ell$-cohesive group with every $\ell\in[k]$, there is at least one voter which approves at least $\ell$ candidates in $W$.
@@
While there are a wealth of concepts associated with JR, there are instances where comparing the performance of different winning committees becomes impossible. To illustrate this situation, consider a brief example.
@@
  • each of the candidates $c_{4},c_{5},c_{6}$ is approved by voters $1,2,4,5,7,8$.
@@ 1.2 Our Technical Contributions
  • We first show that the algorithm for finding a JR committee proposed by Aziz et al. [2] also provides a $\frac{1}{k}$-approximation to MDJR, and the algorithm for finding an EJR committee proposed by Aziz et al. [3] also provides a $\frac{1}{k+1}$-approximation to MDEJR. On the other hand, since the maximum (E)JR degree for both MDJR and MDEJR is $\lceil\frac{n}{k}\rceil$, the approximation guarantees for the two algorithms above can also be written as $1/\lceil\frac{n}{k}\rceil\approx k/n$. To complement these positive results, we show almost tight inapproximability results. We show that it is NP-hard to approximate MDJR (MDEJR) within a factor of $(k/n)^{1-\epsilon}$ for any $\epsilon>0$. We also show that it is NP-hard to approximate MDJR (MDEJR) within a factor of $(1/k)^{1-\epsilon}$ for any $\epsilon>0$.
@@ 2 Inapproximability of MDJR and MDEJR
In this section, we study the approximability of MDJR and MDEJR. We provide almost tight approximability of both problems, in terms of both $k$ and $k/n$.
As we mentioned before, the maximum degree for both MDJR and MDEJR is $\lceil\frac{n}{k}\rceil$. On the other hand, any winner committee satisfying JR (EJR, resp.) gives a JR degree (EJR degree, resp.) of at least $1$. Therefore, the algorithm for finding a JR (EJR, resp.) committee provides a $1/\lceil\frac{n}{k}\rceil\approx\frac{k}{n}$ approximation to MDJR (MDEJR, resp.). In Sect. 2.1, we will show that the algorithm for finding a JR committee proposed by Aziz et al. [2] also provides a $\frac{1}{k}$-approximation to MDJR, and the algorithm for finding an EJR committee proposed by Aziz et al. [3] also provides a $\frac{1}{k+1}$-approximation to MDEJR.
In Sect. 2.2, we show that the approximation ratios of approximately $k/n$ and $1/k$ mentioned above are almost tight.
2.1 Approximation Algorithms
The algorithm GreedyAV (Algorithm 1) proposed by Aziz et al. [2] always outputs a JR committee. We will show in Theorem 1 that it provides a $\frac{1}{k}$-approximation to MDJR.
Input: An instance $\mathcal{I}=(N,C,\mathbf{A},k))$
Output: A winning committee $W$ of size $k$
1 $W\leftarrow\emptyset$
2for $j\in[k]$ do
3       Let $c$ be the candidate approved by the maximum number of voters in $N$
4      $W\leftarrow W\cup c$
5      $N\leftarrow N\setminus V$ where $V$ is the set of voters who approve $c$
return $W$
Algorithm 1 Greedy Approval Voting (GreedyAV)
We prove the following proposition first.
Proposition 2.
GreedyAV outputs a committee achieving JR degree at least $\frac{n}{k^{2}}$.
Proof.
We prove it by contradiction. Suppose the committee output by GreedyAV provides the JR degree less than $\frac{n}{k^{2}}$. Hence, there will be at least $(\frac{n}{k}-\frac{n}{k^{2}}+1)$ voters approving no candidate in $W$ that approve a common candidate $c\notin W$. From the definition of GreedyAV, the coverage of voters is at least $(\frac{n}{k}-\frac{n}{k^{2}}+1)$ in each iteration, and at least $\frac{n}{k}$ in the first iteration (given that there is at least one cohesive group). Therefore, the total number of voters is at least
$\frac{n}{k}+(k-1)\left(\frac{n}{k}-\frac{n}{k^{2}}+1\right)+\left(\frac{n}{k}-\frac{n}{k^{2}}+1\right)=n+\frac{n}{k}-k\left(\frac{n}{k^{2}}-1\right)>n,$
which leads to a contradiction in the number of voters. ∎
Now, we are ready to show the approximation guarantee of GreedyAV.
Theorem 1.
GreedyAV runs in polynomial time and provides a $\frac{1}{k}$-approximation to MDJR.
Proof.
It is clear that the algorithm runs in polynomial time. To show the approximation guarantee, we show that the minimum JR degree of $\frac{n}{k^{2}}$ proved in Proposition 2. To see this, notice that we have remarked that the maximum possible JR degree is $\lceil\frac{n}{k}\rceil$. On the other hand, since the JR degree is an integer, the minimum JR degree that Greedy AV can guarantee is actually $\lceil\frac{n}{k^{2}}\rceil$. It suffices show that
$\left\lceil\frac{n}{k^{2}}\right\rceil\geq\frac{1}{k}\cdot\left\lceil\frac{n}{k}\right\rceil.$
The inequality clearly holds if $n$ is a multiple of $k$. Suppose otherwise. Let $n=ak+b$ for non-negative integers $a,b$ with $11\leq b\leq k-$ 1. Let $a=sk+t$ for non-negative integers $s,t$ with $10\leq t\leq k-$ 1. For the left-hand side of the inequality, we have
$\left\lceil\frac{n}{k^{2}}\right\rceil=\left\lceil\frac{(sk+t)k+b}{k^{2}}\right\rceil\geq\left\lceil s+\frac{b}{k^{2}}\right\rceil\geq s+1,$
where the first inequality above is due to $t\geq 0$ and the second inequality above is due to $b\geq 1$. For the right-hand side of the inequality, we have
$\frac{1}{k}\cdot\left\lceil\frac{n}{k}\right\rceil=\frac{1}{k}\cdot(a+1)=\frac{sk+t+1}{k}\leq s+1,$
where the last inequality is due to that $t\leq k-1$.
Putting together, we have $\left\lceil\frac{n}{k^{2}}\right\rceil\geq\frac{1}{k}\cdot\left\lceil\frac{n}{k}\right\rceil$, which implies the theorem. ∎
$s_{\text{PAV}}(W)=\sum_{i=1}^{n}\sum_{j=1}^{|A_{i}\cap W|}\frac{1}{j}.$
Input: An instance $\mathcal{I}=(N,C,\mathbf{A},k))$
Output: A winning committee $W$ of size $k$
1 $W\leftarrow k$ arbitrary candidates from $C$
3       $W\leftarrow W\setminus\{c^{-}\}\cup\{c^{+}\}$
return $W$
Algorithm 2 $\lambda$-LS-PAV
Next, we will show that $\lambda$-LS-PAV runs in polynomial time and provides a $\frac{1}{k+1}$-approximation to MDEJR for a suitable choice of $\lambda$.
We show the following proposition first.
Proposition 3.
$\lambda$-LS-PAV outputs a committee achieving EJR degree at least $c^{\ast}=\frac{n}{k(k+1)}-\lambda\frac{k}{k+1}$ for any $\lambda\in[0,\frac{n}{k^{2}})$.
Proof.
We prove it by contradiction. Suppose the committee, $W$, output by $\lambda$-LS-PAV provides the EJR degree strictly smaller than $c^{\ast}$. There exists a $\ell$-cohesive group $V\subseteq N$, such that less than $c^{\ast}$ voters in $V$ approves at least $\ell$ members of $W$, i.e., |$|\{i\in V:|A_{i}\cap W|\geq\ell\}|<c^{\ast}$. Since $V$ is $\ell$-cohesive, there exist $\ell$ candidates approved by all voters in $V$. At least one such candidate, $c^{+}\in\bigcap_{i\in V}A_{i}$, is not selected, as otherwise all voters in $V$ approve at least $\ell$ members of $W$.
We will show that there exists a candidate $c^{-}\in W$ such that the increment of the score by swapping $c^{+}$ and $c^{-}$ is at least $\lambda$, so $W$ cannot be an output of $\lambda$-LS-PAV, leading to a contradiction. To see this, we try to swap $c^{+}$ and any candidate $c^{-}\in W$. Since $c^{+}\in A_{i}$ for all $i\in V$, we have
$\Delta(W,c^{+},c^{-})=$ $\sum_{\begin{subarray}{c}i:c^{+}\in A_{i}\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{\begin{subarray}{c}i:c^{+}\notin A_{i}\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
and, by summing up $\Delta(W,c^{+},c^{-})$ for $c^{-}\in W$,
$\sum_{c^{-}\in W}\Delta(W,c^{+},c^{-})\geq$ $\sum_{c^{-}\in W}\sum_{\begin{subarray}{c}i\in V:\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{c^{-}\in W}\sum_{\begin{subarray}{c}i\in N\setminus V:\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
$=$ $\left(\sum_{i\in V}\frac{k-|A_{i}\cap W|}{|A_{i}\cap W|+1}\right)-\left(n-|V|\right)$
$=$ $\left(\sum_{i\in V}\frac{k+1}{|A_{i}\cap W|+1}\right)-n$
$\geq$ $\left(\sum_{i\in V:|A_{i}\cap W|<\ell}\frac{k+1}{|A_{i}\cap W|+1}\right)-n$
$\geq$ $\left(\sum_{i\in V:|A_{i}\cap W|<\ell}\frac{k+1}{\ell}\right)-n.$
Since less than $c^{\ast}$ voters in $V$ approve at least $\ell$ members of $W$, the number of voters in $V$ approve less than $\ell$ members of $W$ is at least $|V|-c^{\ast}+1$. Thus,
$\sum_{c^{-}\in W}\Delta(W,c^{+},c^{-})$ $\geq\left(\sum_{i\in V:|A_{i}\cap W|<\ell}\frac{k+1}{\ell}\right)-n$
$\geq(k+1)\frac{|V|-c^{\ast}+1}{\ell}-n$
$>(k+1)\left(\frac{|V|}{\ell}-c^{\ast}\right)-n$
$\geq(k+1)\left(\frac{n}{k}-c^{\ast}\right)-n$
$=\frac{n}{k}-(k+1)c^{\ast}$
$=\lambda k.$
From the pigeonhole principle, it follows that there exists a candidate $c^{-}\in W$ such that $\Delta(W,c^{+},c^{-})\geq\lambda$. ∎
Now, we are ready to conclude the approximation guarantee of $\lambda$-LS-PAV.
Theorem 2.
For $\lambda=\frac{1}{2k^{2}}$, $\lambda$-LS-PAV runs in polynomial time and provides a $\frac{1}{k+1}$-approximation to MDEJR.
Proof.
Firstly, the algorithm runs in polynomial time: the maximum possible PAV-score is $n\cdot\sum_{j=1}^{k}\frac{1}{j}=O(n\log k)$. On the other hand, the PAV-score is increased by at least $\frac{1}{2k^{2}}$. The algorithm terminates in at most $O(nk^{2}\log k)$ iterations, which runs in polynomial time.
To prove the approximation guarantee, Proposition 3 shows that the EJR degree of the output committee is at least $\frac{n}{k(k+1)}-\frac{1}{2k(k+1)}$. Since the degree is an integer, this is at least $\lceil\frac{n}{k(k+1)}\rceil$: if $\frac{n}{k(k+1)}$ is an integer, we have $\lceil\frac{n}{k(k+1)}-\frac{1}{2k(k+1)}\rceil=\frac{n}{k(k+1)}$ as $\frac{1}{2k(k+1)}<1$; otherwise, since $\frac{n}{k(k+1)}-\lfloor\frac{n}{k(k+1)}\rfloor\geq\frac{1}{k(k+1)}>\frac{1}{2k(k+1)}$, the numbers $\frac{n}{k(k+1)}$ and $\frac{n}{k(k+1)}-\frac{1}{2k(k+1)}$ have the same integral part.
Following a similar analysis as in the proof of Theorem 1, we can prove that
$\left\lceil\frac{n}{k(k+1)}\right\rceil\geq\frac{1}{k+1}\cdot\left\lceil\frac{n}{k}\right\rceil,$
which implies the theorem.
Specifically, the inequality clearly holds if $k$ divides $n$. Otherwise, let $n=ak+b$ for integers $a\geq 0$ and $11\leq b\leq k-$ 1. Let $a=s(k+1)+t$ for integers $s\geq 0$ and $0\leq t\leq k$. We have
$\left\lceil\frac{n}{k(k+1)}\right\rceil=\left\lceil\frac{sk(k+1)+tk+b}{k(k+1)}\right\rceil\geq s+1\geq\frac{s(k+1)+t+1}{k+1}=\frac{1}{k+1}\cdot(a+1)=\frac{1}{k+1}\cdot\left\lceil\frac{n}{k}\right\rceil.\qed$
2.2 Matching Inapproximability Results
To complement the positive results mentioned in the previous section, we present the following inapproximability results.
@@ 2.2 Matching Inapproximability Results
Theorem 5.
It is NP-hard to approximate MDJR within a factor of $\left(\frac{1}{k}\right)^{1-\epsilon}$ for any $\epsilon>0$.
Theorem 6.
It is NP-hard to approximate MDEJR within a factor of $\left(\frac{1}{k}\right)^{1-\epsilon}$ for any $\epsilon>0$.
We will simultaneously prove these four theorems by constructing a hard ballot instance that is used for all of them. We will make sure the instance we constructed has no $\ell$-cohesive group with $\ell>1$. Notice that the JR degree and the EJR degree for any committee are always the same for instances with only $1$-cohesive groups. In addition, we have $n\approx k^{2}$ in our construction, so the factor $k/n$ is approximately $1/k$.
Before we prove the theorems, we first introduce a NP-hard problem: sparse-SAT problem. One can find variations of SAT problem and a similar argument in Tovey [36].
@@
Given a CNF formula $\phi$ that, for any pair of variables $x$ and $y$, at most one clause contains both $x$ (or $\neg x$) and $y$ (or $\neg y$), decide if there is a value assignment to the variables to make $\phi$ true.
To see its NP-hardness, it can be reduced from the SAT problem. Start with any SAT instance. Without loss of generality, suppose each variable, $x$ or $\neg x$, appears in each clause at most once. For each variable $x$ such that $x$ or $\neg x$ is contained in more than two clauses, we perform the following procedure. Suppose $x$ appears in $k$ clauses. Create $k$ new variables $x_{1},\dots,x_{k}$ and replace the $i$th occurrence of $x$ with $x_{i}$ (and $\neg x$ is replaced by $\neg x_{i}$, respectively) for each $i=1,\dots,k$. Add the clause $(x_{i}\vee\neg x_{i+1})$ for $i=1,\dots,k-1$ and the clause $(x_{k}\vee\neg x_{1})$. Note that, in the new instance, variable $x_{i}$ and $y_{j}$ appear in a clause only when the $i$th occurrence of $x$ and the $j$th occurrence of $y$ in $\phi$ are in the same clause, so the new instance satisfies the requirement of the sparse-SAT problem.
@@
[…]
Figure 1: An example of the construction with $\phi=(x_{1}\vee\neg x_{3})\wedge(\neg x_{1}\vee x_{2})$. Edges in the graph represent approvals. An edge connecting between a group of voters and a group of candidates indicates that every voter in the voter group approves every candidate in the candidate group.
Proof of Theorem 3 and 4.
@@
We reduce from sparse-SAT problem. Given any sparse-SAT instance $\phi$, suppose there are $\bar{n}$ clauses and $\bar{m}$ variables (say $x_{1},\dots,x_{\bar{m}}$). We consider an ABC voting instance with $13\bar{m}+\bar{n}+$ 1 candidates $\{c_{1},\dots,c_{2\bar{m}}$, $s_{1},\dots,s_{\bar{n}}$, $t_{1},\dots,t_{\bar{m}},d\}$ and $\bar{n}\bar{m}+\bar{m}^{2}+\bar{n}+\bar{m}$ voters. We want to select a committee of size $\bar{m}+1$. Hence, we care about the cohesive group of size $\frac{\bar{n}\bar{m}+\bar{m}^{2}+\bar{n}+\bar{m}}{\bar{m}+1}=\bar{n}+\bar{m}$. First, for each variable $x_{j}$ and its negation $\neg x_{j}$, we create two corresponding candidates $c_{2j-1}$ and $c_{2j}$ and a group $T_{j}$ of $\bar{m}$ voters who approve $c_{2j-1}$ and $c_{2j}$. For the $i$th clause, we create a group $S_{i}$ of $\bar{m}$ voters. All voters in group $S_{i}$ approve $c_{2j-1}$ if $x_{j}$ occurs in the $i$th clause and $c_{2j}$ if $\neg x_{j}$ occurs in the $i$th clause. All voters in group $S_{i}$ approve $s_{i}$ and voters in group $T_{i}$ approve $t_{i}$ additionally. We create a set $D$ of $\bar{n}$ voters who approve $s_{1},\dots,s_{\bar{n}},t_{1},\dots,t_{\bar{m}},d$. Hence, for each $i\in[\bar{n}]$, $S_{i}\cup D$ forms a $1$-cohesive group, and for each $j\in[\bar{m}]$, $T_{j}\cup D$ forms a $1$-cohesive group. Moreover, we create a set $D^{+}$ of $\bar{m}$ voters who approve $d$. Hence, $D\cup D^{+}$ forms a $1$-cohesive group. An example of our construction is shown in Fig. 1.
@@
Notably, there is no $2$-cohesive group. First, candidate $s_{i}$ for $i\in[\bar{n}]$ or $t_{j}$ for $j\in[\bar{m}]$ or $d$ has only $\bar{n}+\bar{m}$ voters approving them as constructed above. For candidates $c_{1},\dots,c_{2\bar{m}}$, we will show that no $2(\bar{n}+\bar{m})$ voters have two common approved candidates. For any two candidates $t_{2i-1},t_{2i}$ in $\{c_{1},\dots,c_{2\bar{m}}\}$ that correspond to a variable and its negation, the set of voters approving both candidates is exactly $T_{i}$ (notice that we can assume without loss of generality that $x_{i}$ and $\neg x_{i}$ do not appear in the same clause, for otherwise, we can safely remove both literals). We have $|T_{i}|=\bar{m}<2(\bar{n}+\bar{m})$. For any two candidates in $\{c_{1},\dots,c_{2\bar{m}}\}$ that correspond to different variables, by our sparsity assumption of $\phi$, the set of voters approving both candidates is at most some $S_{j}$ (representing a clause that contains both variables). We have $|S_{j}|=\bar{m}<2(\bar{n}+\bar{m})$. Thus, there is no $2$-cohesive group or $\ell$-cohesive group for $\ell>1$. In this case, the EJR degree is equal to the JR degree. Thereafter, we will analyze the JR degree and MDJR, and the analysis applies to MDEJR as well.
@@
If there is a value assignment to the variables to make $\phi$ true, then MDJR will achieve a JR degree of $\bar{n}+\bar{m}$. To see this, for each $j\in[\bar{m}]$, we select $c_{2j-1}$ as a winner if $x_{j}$ is assigned true, or we select $c_{2j}$. Then we select $d$ as the winner. We can verify that each voter in $\{S_{i}\}_{i\in[\bar{n}]}\cup\{T_{j}\}_{j\in[\bar{m}]}$ approves at least one winner. In addition, $d$ is approved by all voters in $D$ and $D^{+}$. Hence, all voters approve at least one winner, and the JR degree equals the size of a $1$-cohesive group, which is $\bar{n}+\bar{m}$.
@@
If there does not exist a satisfying assignment to $\phi$, then MDJR will achieve the JR degree of at most $\bar{n}$. To see this, we prove it by contradiction. Assume that MDJR can achieve the JR degree of larger than $\bar{n}$ by the winner committee $W$ with $|W|=k=\bar{m}+1$. If $d\notin W$, no voter in $D^{+}$ can be covered since they only approve $d$. Hence, for the cohesive group $D\cup D^{+}$, at most $\bar{n}$ voters can be covered, which leads to a contradiction. Thus, $d\in W$. Let $W^{\prime}=W\setminus\{d\}$, and we have $|W^{\prime}|=\bar{m}$. If voters in $T_{j}$ are not covered for some $j\in[\bar{m}]$, at most $\bar{n}$ voters can be covered in the cohesive group $D\cup T_{j}$, leading to a contradiction. Thus, all voters in group $T_{j}$ must be covered for each $j\in[\bar{m}]$, indicating that exactly one of the $3$ candidates $\{c_{2j-1},c_{2j},t_{j}\}$ for each $j\in[\bar{m}]$ is selected (at least one candidate in each three-candidates group must be selected, we have $\bar{m}$ groups, and we have $|W^{\prime}|=\bar{m}$). Next, we will show that voters in at least one group among $S_{1},\dots,S_{\bar{n}}$ cannot be covered. Suppose this is not the case. For every $t_{j}\in W^{\prime}$, we can use either $c_{2j-1}$ or $c_{2j}$ to replace $t_{j}$ since $t_{j}$ only covers group $T_{j}$ and each of $c_{2j-1}$ and $c_{2j}$ covers at least $T_{j}$. Hence, we can find one candidate in $\{c_{2i-1},c_{2i}\}$ for all $i\in[\bar{m}]$ to cover all groups, implying that there exists a value assignment to the variables ($x_{i}$ is assigned true if $c_{2i-1}$ is selected, or false otherwise) to make $\phi$ true, leading to a contradiction. Therefore, at least one group among $S_{1},\dots,S_{\bar{n}}$ cannot be covered. Without loss of generality, we assume that it is $S_{i}$. Then, for the cohesive group $S_{i}\cup D$, at most $\bar{n}$ voters can be covered, leading to a contradiction.